OpenStudy (anonymous):

What is the volume of this triangular prism? A. 216 cm3 B. 432 cm3 C. 864 cm3 D. 38 cm3

OpenStudy (_michaeld22):

Can you please post the diagram of it?

OpenStudy (anonymous):



OpenStudy (anonymous):

it lok\oks like crap because im 3D blind and i cant see 3D

OpenStudy (anonymous):

*looks

OpenStudy (_michaeld22):

\[V=\frac{ 1 }{ 2 }(l x w x h)\] 12*8 = 96. 96*18 = 1728. 1728/2 = 864.

OpenStudy (_michaeld22):

Sorry those x's are supposed to be multiplication signs >.<